OpenStudy (anonymous):

A rectangular box is x feet long and x feet wide. The volume of the box is (4x^8+3x^6) cubic feet. What polynomial represents the height of the box in feet? A.4x+3 B.4x^6+3x^4 C.4x^8+3x^6 D.4x^7+3x^5

OpenStudy (anonymous):

The answer should be B. Volume=width*height*length so V=x*x*h=4x^8+3x^6 x^2*h=x^4(4x^2+3) isolate h on the left side to solve for it h=x^4(4x^2+3)/x^2 h=4x^6+3x^4

OpenStudy (anonymous):

thank you